Suppose an investment offers to triple your money in 27 years. What annual rate of return are you being offered if interest is compounded monthly?

Group of answer choices

1.42

5.62

2

4.08
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value($3x)
P=present value($x say)
r=rate of interest
n=time period.

3x=x*(1+r/12)^(12*27)

(3)^(1/324)=(1+r/12)

(1+r/12)=1.003396534

r=(1.003396534-1)*12

=4.08%(Approx).